OpenStudy (anonymous):

There are two solutions being mixed by a chemist. The first consists of 65% chlorine and the second has 40% chlorine. How many liters of the second solution will be needed for a 25–liter mixture that is 57% chlorine?

OpenStudy (anonymous):

put \(.65x+.4(25-x)=25\times .57\) and solve for \(x\)

OpenStudy (anonymous):

probably easier to multiply by 100 and solve \[65x+40(25-x)=25\times 57\]